# Shafarevich’s conjecture for families of hypersurfaces over function fields

### v1

@@ 4. Examples of Infinitesimal Torelli
For moduli space of cubic surfaces, even though the period map of relative middle cohomology is trivial, the occult period map constructed in [ACT02] provides a period map which satisfies infinitesimal Torelli.
@@
By Proposition 4.1 and Proposition 4.2, we can take ${\mathcal{M}}={\mathcal{U}}_{n,d}$ and ${\mathcal{M}}={\mathcal{I}}_{\underline{d}}$ in Theorem 1.8. By Beauville [Bea86, Théorèmes 2,4,5,6], the generic Mumford–Tate group is ${\mathbb{Q}}$-simple and acts absolutely irreducibly. For cubic surfaces, we use the period map from [ACT02] which satisfies infinitesimal Torelli and it results from [ACT02, Theoerm 2.14] that the generic Mumford–Tate group of the family acts absolutely irreducibly. This proves Theorem 1.4 and Theorem 1.7. ∎

### v2

@@ 4. Examples of Infinitesimal Torelli
For the moduli space of cubic surfaces, even though the period map of relative middle cohomology is trivial, the occult period map constructed in [ACT02] provides a period map which satisfies infinitesimal Torelli.
@@
By Proposition 4.1 and Proposition 4.2, we can take ${\mathcal{M}}={\mathcal{U}}_{n,d}$ and ${\mathcal{M}}={\mathcal{I}}_{\underline{d}}$ in Theorem 1.9. By Beauville [Bea86, Théorèmes 2,4,5,6] [^12] See also [PS03, Corollary 18]., the generic Mumford–Tate group is ${\mathbb{Q}}$-simple, therefore its derived group is equal to the algebraic monodromy and acts absolutely irreducibly. For cubic surfaces, we use the period map from [ACT02] which satisfies infinitesimal Torelli and it results from [ACT02, Theorem 2.14] that the generic Mumford–Tate group of the family is equal to its monodromy. This proves Theorem 1.4 and Theorem 1.7. ∎